# Pentagonal nanowires from topological crystalline insulators: a platform for intrinsic core-shell nanowires and higher-order topology

### v1

@@ Abstract
We report on the experimental realization of PbSnTe pentagonal nanowires (NWs) with [110] orientation using molecular beam epitaxy techniques. Using first-principles calculations, we investigate the structural stability in NWs of SnTe and PbTe in three different structural phases: cubic, pentagonal with [001] orientation and pentagonal with [110] orientation. Within a semiclassical approach, we show that the interplay between ionic and covalent bonds favors the formation of pentagonal NWs. Additionally, we find that this pentagonal structure is more likely to occur in tellurides than in selenides. The disclination and twin boundary cause the electronic states originating from the NW core region to generate a conducting band connecting the valence and conduction bands, creating a symmetry-enforced metallic phase. The metallic core band has opposite slopes in the cases of Sn and Te twin boundary, while the bands from the shell are insulating. We finally study the electronic and topological properties of pentagonal NWs unveiling their potential as a new platform for higher-order topology and fractional charge. These pentagonal NWs represent a unique case of intrinsic core-shell one-dimensional nanostructures with distinct structural, electronic and topological properties between the core and the shell region.
@@ I Introduction
Very few real materials have been proposed for the experimental realization of HOTI, two of the most popular are bismuthSchindler et al. (2018a) and cubic SnTe nanowires (NWs)Schindler et al. (2018b). Indeed, it was proposed that the cubic SnTe NWs under rhombohedral strain can host a helical higher-order topological insulator phase Schindler et al. (2018b). This is protected by both mirror and time-reversal symmetry, but no inversion symmetry is needed for helical HOTI in SnTe. The rhombohedral strain has the role of suppressing the first-order topology to make the second-order topology visibleSchindler et al. (2018b). Recently, it was proposed a new route through defects or dislocations to obtain higher-order topological insulatorsGeier et al. (2021); Schindler et al. (2022); Benalcazar et al. (2014a); Roy and Juričić (2021a) or fractional chargesPeterson et al. (2021); Liu et al. (2021a). In case of disclinations, the system can host
@@
strong second-order topological phases provided that local C, C, or C symmetries protecting the topological phase are presentGeier et al. (2021); the same happens for one-dimensional disclinationsZhang et al. (2023).
@@
Group IV-VI semiconductors are known to display interesting characteristics, including thermoelectric effects Wang et al. (2011); Wood (1988), ferroelectric behavior Lebedev and Sluchinskaya (1994); Liu et al. (2021b), and superconducting properties Matsushita et al. (2006); Mazur et al. (2019). Notably, the identification of a topological crystalline insulating phase (TCI) in materials such as SnTe Fu (2011); Hsieh et al. (2012); Lau and Ortix (2019); Tanaka et al. (2012); Okada et al. (2013), along with certain related substitutional compounds such as PbSnTe Xu et al. (2012) and PbSnSe Dziawa et al. (2012); Łusakowski et al. (2023), has sparked significant interest in the realm of research. This category of materials has garnered substantial attention for further investigation. The TCI phase stands apart from the conventional topological insulating phase, as the linearly dispersing Dirac states on the high-symmetry surfaces of TCIs are protected by crystal symmetriesFu et al. (2007) rather than time-reversal symmetry. Consequently, the topological invariant is also different. For example, in the case of SnTe and its alloys, the invariant is named mirror Chern number Teo et al. (2008); Safaei et al. (2015); Hsieh et al. (2012); Cuono et al. (2022); Rechciński et al. (2021) since the symmetry which protects the surface states is a reflection of the crystal with respect to (110) mirror plane. Changing the dimensions or the size (whether in a 2D or 3D configuration) of materials results in the modification of their distinctive properties. Notably, a transition from bulk Plekhanov et al. (2014); Wang et al. (2020); Barone et al. (2013) to thin films Sławińska et al. (2020); Volobuev et al. (2017); Liu et al. (2018) induces property variations. A case in point is SnTe: while it behaves as a trivial insulator at reduced thicknesses, it assumes a topological nature beyond a critical thicknessLiu et al. (2014, 2015); Ozawa et al. (2014); Safaei et al. (2015); Liu and Fu (2015), displaying strong robustness against impurity dopingArroyo-Gascón et al. (2022). Employing scanning tunneling microscopy and spectroscopy on PbSnSe, it was demonstrated that the (001) surface hosts large terraces interspersed with step edges of varying heights. The features of low-energy states occurring at surface atomic steps have been investigated, along with an exploration of the response of these edge channels to the stoichiometry of the alloySessi et al. (2016); Brzezicki et al. (2019); Mazur et al. (2019); Wagner et al. (2023); Rechciński and Buczko (2018).
@@
The effects of the presence of defects in the TCIs were also investigated in the literature. Possible defects are disclinations or dislocations, which naturally occur in the material during its growth. These are defects that interrupt the rotational or translational symmetry in a region of the material. Fractional electric charges were observed in disclinations Ortix (2016); Peterson et al. (2021); Liu et al. (2021c), it was shown that these defects can probe higher-order topology Schindler et al. (2018b, a); Roy and Juričić (2021b), and the criteria for a zero-energy Majorana bound state to be present at defects was given Benalcazar et al. (2014b); Chen et al. (2022). Partial-dislocation-induced topological modes in 2D and 3D higher-order topological insulators were experimentally observed Yamada et al. (2022). In the recent past, an investigation was undertaken to explore the connection between the twinning in thin films and variations in the mirror Chern numberSamadi et al. (2023).
@@
Exploring the one-dimensional (1D) case, the topological crystalline insulator SnTe NWs (NWs) emerge as even more captivating than their bulk counterparts. This enhanced intrigue arises from the amplified contribution originating from topological surface states and the effects of confinement in 1D systemsShen and Cha (2014); Xu et al. (2015). Diverse approaches were uncovered to produce SnTe NWs. These methods encompass the utilization of high-yield alloy nanoparticles as growth catalysts Liu et al. (2021b), graphene Sadowski et al. (2018) and employing single-walled carbon nanotubes to produce the smallest possible SnTe NWs Vasylenko et al. (2018). SnTe NWs have also been thoroughly investigated theoretically. It was shown that it is possible to obtain distinct topological states through the application of varied fields Nguyen et al. (2022). Notably, bulk Majorana modes appear in the presence of inversion symmetry, while the imposition of symmetry-breaking fields at the NW ends leads to the emergence of Majorana zero modes, alongside the gapping of the Majorana bulk modes Nguyen et al. (2022). Cubic PbSnTe NWs have also been synthesized Safdar et al. (2015); Dad et al. (2024); Saghir et al. (2015); Xu et al. (2016). TCI NWs can serve as a platform to confine and control Dirac fermions Skiff et al. (2023), while PbTe NWs have drown attention due to their potential to host Majorana zero modes Song et al. (2023); Jung et al. (2022). However, the realization of Majorana’s fermions is challenging as it requires the growth of NWs to be integrated with a circuit that exhibits the necessary sensitivity to monitor individual electrons as they traverse through it. All of this intricate work must be conducted at extremely low temperatures and within a magnetic field that is 10,000 times stronger than Earth’sFrolov (2021).
@@
Ab initio calculations were applied to investigate ultrathin cubic NWsShukla et al. (2022); Hussain et al. (2023) and to find the critical thicknesses at which SnTe NWs go from a trivial insulating regime to a spin-orbit insulating phase and then to the TCI phase Hussain et al. (2023). Great attention was recently dedicated to 2D pentagonal materials in different types of the Cairo pentagonal tessellationShen and Wang (2022); Winczewski and Rybicki (2022); Bykov et al. (2021) and their topological propertiesBravo et al. (2019). Studies are present also for 1D pentagonal materials; however, only elemental and non-ionic pentagonal NWs have been experimentally realizedGonzález et al. (2004); Reyes-Gasga et al. (2006); Antsov et al. (2019) or theoretically proposedSen et al. (2002); Ma and Zhao (2013); Sainath and Choudhary (2016); Pahl et al. (2008).
@@
First, we report the experimental realization of PbSnTe ionic pentagonal NWs using molecular beam epitaxy technique. Then we investigate, by means of ab initio and model Hamiltonian calculations, the structural, electronic and topological properties of pentagonal PbTe and SnTe NWs with anionic and cationic twin boundaries.
We find that the cubic [001] NWs show a robust insulating behavior Hussain et al. (2023), while the pentagonal phases are metastable with a more covalent bond. This more covalent bond brings smaller band gaps and therefore favors the topological phases.
@@ II Experimental realization of ionic pentagonal nanowires
PbSnTe NWs were grown using the molecular beam epitaxy (MBE) approach within a home-built apparatus furnished with solid sources of SnTe, Pb, and Te (refer to Appendix A for more comprehensive information). The NWs obtained with this methods have typical lengths exceeding 1 m, with a NW growth rate surpassing 0.5 m/h. The NWs diameters display a relatively broad range, spanning from a minimum of approximately 15 nm to as much as 200 nm. The peak distribution rests at around 40-80 nm. Under these specific conditions, both varieties of NWs are observed: those manifesting a four-fold and those with a five-fold symmetry in their cross-sectional profiles (see Fig. 1a).
Utilizing a Scanning Electron Microscope (SEM), we conducted observations at 5 kV, with magnifications ranging from 10k to 300k x. The SEM imagery reveals a diverse NW length, spanning from 0.7 m to 3 m, while the width of the structures spans between 40 nm and 150 nm. On the substrate, a prevalent occurrence of five-fold NWs was noted, with less frequent instances of four-fold NWs. These structures exhibit remarkable symmetry and well-defined shapes, displaying evident blocks. The NW surfaces exhibit a flawlessly smooth texture devoid of imperfections.
To explore the structural attributes of the PbSnTe NWs, we employed a transmission electron microscope. This analysis involved cross-sectioned NWs inspected along the [110] zone axis, which aligns with the NW growth axis. Both scanning transmission electron microscopy and high-resolution transmission electron microscopy modes were employed. The presence of five-fold symmetry was conclusively established. Furthermore, distinct features such as the (111) plane twin boundaries were discernible, along with brighter areas indicative of twinning. These areas are related to strain fields. In Fig. 1b), we show the SEM top view image of PbSnTe NWs and the SEM image of five-fold PbSnTe NW. At the core of the pentagonal structure, we have a one-dimensional disclination that follows the length of the pentagonal NW. The disclination is described by the Frank anglePeterson et al. (2021), which in this case is +90. More technical details are reported in Appendix A. The twin boundaries of the experimentally obtained pentagonal nanowires are composed of cation atoms. This will be shown somewhere else, while the next theoretical Sections will investigate both cation and anion twin boundaries.
@@ III.1 Structural stability of the nanowires
Using density functional theory (DFT) calculations, we compare the energetic stability of three NW crystal structures. The computational details of the calculations are reported in Appendix B. We study the cubic NWs with [001] growth orientation and the two growth orientations [001] and [110] for the pentagonal NWs. The pentagonal phases have 5 mirror planes that coincide with the plane of the twin boundaries. The pentagonal phase with [110] orientation is the phase observed in experiments. The three optimized structures of thin SnTe NWs are illustrated in Fig. 2. In Fig. 3, we report the side and top view of the [110] pentagonal NWs. The [110] pentagonal NW is composed of a core region and a shell region. The core region is composed of a core chain (CC) along the z-axis where the atoms see an environment with rotational symmetry C. Additionally, there is a mirror plane at z=0 or z=0.5, but there is no inversion point. The forces acting on the atoms of the core chain (one Sn and one Te atom per unit cell) are zero by symmetry. Indeed, the forces along the z-axis are compensated by the mirror symmetry while for the in-plane components, the sum of the forces of equivalent atoms located at the vertex of a regular pentagon on the center is zero. The shell region is composed of the 5 twin boundaries that separate regions where the system is locally cubic and where the atoms see locally the C symmetry present in the bulk system. As we can see, the thickness of these NWs can be increased by adding more rings of corresponding atoms to the edges (sides) of each NW. In this way, the pentagonal NWs are named as 1-ring, 2-ring, 3-ring, 4-ring and 5-ring, respectively.
@@
Here, and are the numbers of cation atoms (Sn, Pb) and Te atoms in the NW, respectively. E and E represent the ground state total energies for individual atoms (E, E, E), and the term E represents the total energy of the relaxed NW i.e. E or E. The higher the binding energy, the higher the structural stability. Fig. 4 shows the binding energy (E) for SnTe and PbTe NWs oriented along [001] and [110] directions, as a function of NW thickness. Also, E of the bulk system is portrayed as a line for comparison. It can be seen that E varies with the number of atoms in the NW. As the thickness of the NW increases, the binding energy also increases. This monotonic increase of E reflects enhancement in the stability of NWs. From Fig. 4, we observed that pentagonal NWs grown in [001] direction bear more stability and are closer to bulk value than NWs grown along [110] orientation. However, the [001] cubic phase of NWs is observed to be the most stable one as compared to other cases. Following this analysis, the cubic phase is still the ground state for SnTe and PbTe NWs as in the bulk phase, while the pentagonal structure of SnTe and PbTe NWs is a metastable state. We calculate the binding energy as a function of the number of atoms in the unit cell, and in the limit of infinite atoms, the binding energy should tend to the value of the bulk. Indeed, the binding energy differences strongly decrease as a function of the thicknesses of NWs and could reduce to a few tens of meV for thick NWs. Therefore, it is unlikely to have thin pentagonal NWs, while they are more likely to occur for large thicknesses. Recently, it was shown that several structural configurations are possible in the case of elemental NWs with a small difference in the binding energy of few meVTarkowski and Gonzalez Szwacki (2023). In the ionic NWs as SnTe, the number of configurations would be certainly smaller with respect to elemental NWs due to the constraints on the alternation of anion and cation; however, as we demonstrated experimentally, it is possible to stabilize structural configurations different from the lowest energy crystal structure.
@@
A semiclassical approach to describing the stability of ionic pentagonal NWs is reported in Appendix C. Within the semiclassical approach, for the cube, the Madelung constant is highest and particle density is lowest, indicating that the contribution of ionic bond is larger than the covalent one. In the case of hexagon, particle density is the highest and Madelung constant is the lowest indicating that the covalent bond is the dominant character of the binding. However, the pentagonal structure bears intermediate values, which implies that if the bond is partially ionic and partially covalent, the pentagonal system is expected to be the most stable one. Basically, the pentagonal structure arises from the interplay between the ionic bond and the covalent bond. The pentagonal phases of the system tend to show a more covalent bond that favors smaller gaps and metallicity. NWs based on Pb and Te would be more likely to form this unusual pentagonal structure since they will have weaker ionic bonds.
@@
Further research is necessary to advance towards the alloyed phase PbSnTe. Our preliminary results indicate that the difference between the binding energy of the cubic and pentagonal NWs decreases to smaller values in the alloy.Hussain and Autieri (2024) Exploring the possibility that the CC consists solely of either Sn or Pb atoms could be a direction worth investigating. In this regard, employing genetic algorithms might predict the crystal structure and the cationic composition of the CCTarkowski and Gonzalez Szwacki (2023).

### v3

@@ Abstract
We report on the experimental realization of Pb1-xSnxTe pentagonal nanowires (NWs) with [110] orientation using molecular beam epitaxy techniques. Using first-principles calculations, we investigate the structural stability in NWs of SnTe and PbTe in three different structural phases: cubic, pentagonal with [001] orientation and pentagonal with [110] orientation. Within a semiclassical approach, we show that the interplay between ionic and covalent bonds favors the formation of pentagonal NWs. Additionally, we find that this pentagonal structure is more likely to occur in tellurides than in selenides. The disclination and twin boundary cause the electronic states originating from the NW core region to generate a conducting band connecting the valence and conduction bands, creating a symmetry-enforced metallic phase. The metallic core band has opposite slopes in the cases of Sn and Te twin boundary, while the bands from the shell are insulating. We finally study the electronic and topological properties of pentagonal NWs unveiling their potential as a new platform for higher-order topology and fractional charge. These pentagonal NWs represent a unique case of intrinsic core-shell one-dimensional nanostructures with distinct structural, electronic and topological properties between the core and the shell region.
@@ I Introduction
Very few real materials have been proposed for the experimental realization of HOTI, two of the most popular are bismuthSchindler et al. (2018) and cubic SnTe nanowires (NWs)Schindler et al. (2018). Indeed, it was proposed that the cubic SnTe NWs under rhombohedral strain can host a helical higher-order topological insulator phase Schindler et al. (2018). This is protected by both mirror and time-reversal symmetry, but no inversion symmetry is needed for helical HOTI in SnTe. The rhombohedral strain has the role of suppressing the first-order topology to make the second-order topology visibleSchindler et al. (2018). Recently, it was proposed a new route through defects or dislocations to obtain higher-order topological insulatorsGeier et al. (2021); Schindler et al. (2022); Benalcazar et al. (2014); Roy and Juričić (2021) or fractional chargesPeterson et al. (2021); Liu et al. (2021). In case of disclinations, the system can host
@@
strong second-order topological phases provided that local C2, C4, or C6 symmetries protecting the topological phase are presentGeier et al. (2021); the same happens for one-dimensional disclinationsZhang et al. (2023).
@@
Group IV-VI semiconductors are known to display interesting characteristics, including thermoelectric effects Wang et al. (2011); Wood (1988), ferroelectric behavior Lebedev and Sluchinskaya (1994); Liu et al. (2021), and superconducting properties Matsushita et al. (2006); Mazur et al. (2019). Notably, the identification of a topological crystalline insulating phase (TCI) in materials such as SnTe Fu (2011); Hsieh et al. (2012); Lau and Ortix (2019); Tanaka et al. (2012); Okada et al. (2013), along with certain related substitutional compounds such as Pb1-xSnxTe Xu et al. (2012) and Pb1-xSnxSe Dziawa et al. (2012); Łusakowski et al. (2023), has sparked significant interest in the realm of research. This category of materials has garnered substantial attention for further investigation. The TCI phase stands apart from the conventional topological insulating phase, as the linearly dispersing Dirac states on the high-symmetry surfaces of TCIs are protected by crystal symmetriesFu et al. (2007) rather than time-reversal symmetry. Consequently, the topological invariant is also different. For example, in the case of SnTe and its alloys, the invariant is named mirror Chern number Teo et al. (2008); Safaei et al. (2015); Hsieh et al. (2012); Cuono et al. (2022); Rechciński et al. (2021) since the symmetry which protects the surface states is a reflection of the crystal with respect to (110) mirror plane. Changing the dimensions or the size (whether in a 2D or 3D configuration) of materials results in the modification of their distinctive properties. Notably, a transition from bulk Plekhanov et al. (2014); Wang et al. (2020); Barone et al. (2013) to thin films Sławińska et al. (2020); Volobuev et al. (2017); Liu et al. (2018) induces property variations. A case in point is SnTe: while it behaves as a trivial insulator at reduced thicknesses, it assumes a topological nature beyond a critical thicknessLiu et al. (2014, 2015); Ozawa et al. (2014); Safaei et al. (2015); Liu and Fu (2015), displaying strong robustness against impurity dopingArroyo-Gascón et al. (2022). Employing scanning tunneling microscopy and spectroscopy on Pb1-xSnxSe, it was demonstrated that the (001) surface hosts large terraces interspersed with step edges of varying heights. The features of low-energy states occurring at surface atomic steps have been investigated, along with an exploration of the response of these edge channels to the stoichiometry of the alloySessi et al. (2016); Brzezicki et al. (2019); Mazur et al. (2019); Wagner et al. (2023); Rechciński and Buczko (2018).
@@
The impacts resulting from the existence of defects in the TCIs were also investigated in the literature. Possible defects are disclinations or dislocations, which naturally occur in the material during its growth. These are defects that interrupt the rotational or translational symmetry in a region of the material. Fractional electric charges were observed in disclinations Ortix (2016); Peterson et al. (2021); Liu et al. (2021), it was shown that these defects can probe higher-order topology Schindler et al. (2018, 2018); Roy and Juričić (2021), and the criteria for a zero-energy Majorana bound state to be present at defects was given Benalcazar et al. (2014); Chen et al. (2022). Partial-dislocation-induced topological modes in 2D and 3D higher-order topological insulators were experimentally observed Yamada et al. (2022). In the recent past, an investigation was undertaken to explore the connection between the twinning in thin films and variations in the mirror Chern numberSamadi et al. (2023).
@@
Exploring the one-dimensional (1D) case, the topological crystalline insulator SnTe NWs emerge as even more captivating than their bulk counterparts. This enhanced intrigue arises from the amplified contribution originating from topological surface states and the effects of confinement in 1D systemsShen and Cha (2014); Xu et al. (2015); Safdar et al. (2013); Li et al. (2013). Diverse approaches were uncovered to produce SnTe NWs. These methods encompass the utilization of high-yield alloy nanoparticles as growth catalysts Liu et al. (2021), graphene Sadowski et al. (2018) and employing single-walled carbon nanotubes to produce the smallest possible SnTe NWs Vasylenko et al. (2018). SnTe NWs have also been thoroughly investigated theoretically. It was shown that it is possible to obtain distinct topological states through the application of varied fields Nguyen et al. (2022). Notably, bulk Majorana modes appear in the presence of inversion symmetry, while the imposition of symmetry-breaking fields at the NW ends leads to the emergence of Majorana zero modes, alongside the gapping of the Majorana bulk modes Nguyen et al. (2022). Cubic Pb1-xSnxTe NWs have also been synthesized Safdar et al. (2015); Dad et al. (2024); Saghir et al. (2015); Xu et al. (2016). TCI NWs can serve as a platform to confine and control Dirac fermions Skiff et al. (2023), while PbTe NWs have drawn attention due to their potential to host Majorana zero modes Song et al. (2023); Jung et al. (2022). However, the realization of Majorana’s fermions is challenging as it requires the growth of NWs to be integrated with a circuit that exhibits the necessary sensitivity to monitor individual electrons as they traverse through it. All of this intricate work must be conducted at extremely low temperatures and within a magnetic field that is 10,000 times stronger than Earth’s fieldFrolov (2021).
@@
Ab initio calculations were applied to investigate ultrathin cubic NWsShukla et al. (2022); Hussain et al. (2024) and to find the critical thicknesses at which SnTe NWs go from a trivial insulating regime to a spin-orbit insulating phase and then to the TCI phase Hussain et al. (2024). Great attention was recently dedicated to 2D pentagonal materials in different types of the Cairo pentagonal tessellationShen and Wang (2022); Winczewski and Rybicki (2022); Bykov et al. (2021) and their topological propertiesBravo et al. (2019). Studies are present also for 1D pentagonal materials referred as penta-twinned nanowires; however, only elemental and non-ionic pentagonal NWs have been experimentally realizedGonzález et al. (2004); Reyes-Gasga et al. (2006); Antsov et al. (2019) or theoretically proposedSen et al. (2002); Ma and Zhao (2013); Sainath and Choudhary (2016); Pahl et al. (2008).
@@
First, we report the experimental realization of Pb1-xSnxTe ionic pentagonal NWs using molecular beam epitaxy technique. Then we investigate, by means of ab initio and model Hamiltonian calculations, the structural, electronic and topological properties of pentagonal PbTe and SnTe NWs with anionic and cationic twin boundaries.
We find that the cubic [001] NWs show a robust insulating behavior Hussain et al. (2024), while the pentagonal NWs are metastable with a more covalent bond. This more covalent bond brings smaller band gaps and therefore favors the topological phases.
@@ II Experimental realization of ionic pentagonal nanowires
Pb1-xSnxTe NWs were grown using the molecular beam epitaxy (MBE) approach within a home-built apparatus furnished with solid sources of SnTe, Pb, and Te (refer to Appendix A in the Supp. Information for more comprehensive information). The NWs obtained with this method have typical lengths exceeding 1 m, with a NW growth rate surpassing 0.5 m/h. The NW diameters display a relatively broad range, spanning from a minimum of approximately 15 nm to as much as 200 nm. The peak distribution rests at around 40-80 nm. Under these specific conditions, both varieties of NWs are observed: those manifesting a four-fold and those with a five-fold symmetry in their cross-sectional profiles (see Fig. 1a). Utilizing a Scanning Electron Microscope (SEM), we conducted observations at 5 kV, with magnifications ranging from 10k to 300k x. The SEM imagery reveals a diverse NW length, spanning from 0.7 m to 3 m, while the width of the structures spans between 40 nm and 150 nm. On the substrate, a prevalent occurrence of five-fold NWs was noted, with less frequent instances of four-fold NWs. These structures exhibit remarkable symmetry and well-defined shapes, displaying evident blocks. The NW surfaces exhibit a flawlessly smooth texture devoid of imperfections.
To explore the structural attributes of the Pb1-xSnxTe NWs, we employed a transmission electron microscope. This analysis involved cross-sectioned NWs inspected along the [110] zone axis, which aligns with the NW growth axis. Both scanning transmission electron microscopy (STEM) and high-resolution transmission electron microscopy modes were employed. The presence of five-fold symmetry was conclusively established. Furthermore, distinct features such as the (111) plane twin boundaries were discernible, along with brighter areas indicative of twinning. These areas are related to strain fields. In Fig. 1b), we show the STEM top view image of the five-fold Pb1-xSnxTe NW. At the core of the pentagonal structure, we have a one-dimensional disclination that follows the length of the pentagonal NW. The disclination is described by the Frank anglePeterson et al. (2021), which in this case is +90∘. Despite having cubic domains between the twin boundaries, the pentagonal NWs present a 1D disclination and five twin-boundary planes; even if these planes are extended defects, their effect on the structure is local so that most of the whole volume of the NW is in cubic arrangement making this phase equivalent to the cubic structure from the thermodynamic point of viewMarks and Peng (2016). From the experimental data, the twin boundaries of the experimentally obtained pentagonal nanowires seem to be composed of anion atoms.
@@ III.1 Structural stability of the nanowires
Using density functional theory (DFT) calculations, we compare the energetic stability of three NW crystal structures. The computational details of the calculations are reported in Appendix B in the Supp. Information. We study the cubic NWs with [001] growth orientation and the two growth orientations [001] and [110] for the pentagonal NWs. The pentagonal structures have 5 mirror planes that coincide with the plane of the twin boundaries. The pentagonal structure with [110] orientation is the phase observed in experiments. The three optimized structures of thin SnTe NWs are illustrated in Fig. 2. In Fig. 3, we report the side and top view of the [110] pentagonal NWs. The [110] pentagonal NW is composed of a core region and a shell region. The core region is composed of a core chain (CC) along the z-axis where the atoms see an environment with rotational symmetry C5. Additionally, there is a mirror plane at z=0 or z=0.5, but there is no inversion point. The forces acting on the atoms of the core chain (one Sn and one Te atom per unit cell) are zero by symmetry. Indeed, the forces along the z-axis are compensated by the mirror symmetry while for the in-plane components, the sum of the forces of equivalent atoms located at the vertex of a regular pentagon on the center is zero. The shell region is composed of the 5 twin boundaries that separate regions where the system is locally cubic and where the atoms see locally the C4 symmetry present in the bulk system. As we can see, the thickness of these NWs can be increased by adding more rings of corresponding atoms to the edges (sides) of each NW. In this way, the pentagonal NWs are named as 1-ring, 2-ring, 3-ring, 4-ring and 5-ring, respectively.
@@
Here, and are the numbers of cation atoms (Sn, Pb) and Te atoms in the NW, respectively. EX and EY represent the ground state total energies for individual atoms (ESn, EPb, ETe), and the term EXY represents the total energy of the relaxed NW i.e. ESnTe or EPbTe. The higher the binding energy, the higher the structural stability. Fig. 4 shows the binding energy (EB) for SnTe and PbTe NWs oriented along [001] and [110] directions, as a function of NW thickness. Also, EB of the bulk system is portrayed as a line for comparison. It can be seen that EB varies with the number of atoms in the NW. As the thickness of the NW increases, the binding energy also increases. This monotonic increase of EB reflects enhancement in the stability of NWs. From Fig. 4, we observed that pentagonal NWs grown in [001] direction bear more stability and are closer to bulk value than NWs grown along [110] orientation. However, the [001] cubic structure of NWs is observed to be the most stable one as compared to other cases. Following this analysis, the cubic structure is still the ground state for SnTe and PbTe NWs as in the bulk phase, while the pentagonal structure of SnTe and PbTe NWs is a metastable state. We calculate the binding energy as a function of the number of atoms in the unit cell, and in the limit of infinite atoms, the binding energy should tend to the value of the bulk. Indeed, the binding energy differences strongly decrease as a function of the thicknesses of NWs and could reduce to a few tens of meV for thick NWs. Therefore, it is unlikely to have thin pentagonal NWs, while they are more likely to occur for large thicknesses. Recently, it was shown that several structural configurations are possible in the case of elemental NWs with a small difference in the binding energy of few meVTarkowski and Gonzalez Szwacki (2023). In the ionic NWs as SnTe, the number of configurations would be certainly smaller with respect to elemental NWs due to the constraints on the alternation of anion and cation; however, as we demonstrated experimentally, it is possible to stabilize structural configurations different from the lowest energy crystal structure. Experimental information supporting the metastable property of the pentagonal NW comes from the growth temperature. As mentioned in the previous Section, the pentagonal NWs are obtained only in high-temperature experiments, where metastable phases are more likely to be obtained.
DFT calculations of the pentagonal nanowires with Pb0.50Sn0.50Te composition do not change the scenario from the point of view of the total energy, indicating that the formation of pentagonal NWs comes from a less likelihood to obtain Pb0.50Sn0.50Te cubic NWs. In order to further understand the stability of the pentagonal nanowires, we discuss what is peculiar in the Pb0.50Sn0.50Te. Well, close to this composition we have the closure of the band gap with the topological transition in the bulk samples or thick cubic NWs at room temperatureDornhaus et al. (1983). The cubic structures of Pb1-xSnxTe are stabilized by the ionic insulator character described by the Madelung constant, while metallic systems tend to have a more hexagonal structure to produce a large atomic packing factor and covalent bonds. We propose that the thick cubic Pb0.50Sn0.50Te NWs with zero gap are structurally less stable than SnTe and PbTe NWs, such that the pentagonal Pb0.50Sn0.50Te NWs are more likely to grow. To test this claim at the quantum mechanical level, we would need to reproduce the energetic stability of the alloyed cubic and pentagonal NWs within DFT, however, we are not able to simulate cubic NWs with spin-orbit above 15 nm thicknessHussain et al. (2024). Therefore, we move to a semiclassical approach to provide insight on the formation of the disclination in the initial stages of the growth.
@@
A semiclassical approach to describing the stability of ionic pentagonal NWs is reported in Appendix C in the Supp. Information. Within the semiclassical approach, for the cube, the Madelung constant is highest and particle density is lowest, indicating that the contribution of the ionic bond is larger than the covalent one. In the case of hexagon, particle density is the highest and Madelung constant is the lowest indicating that the covalent bond is the dominant character of the binding. However, the pentagonal structure bears intermediate values, which implies that if the bond is partially ionic and partially covalent, the pentagonal system is expected to be the most stable one. Basically, the pentagonal structure arises from the interplay between the ionic bond and the covalent bond. The pentagonal NW tends to be stabilized in a more covalent environment which is favored by small gaps and metallicity as happens close to the topological transition in Pb0.50Sn0.50Te. In the Supporting Information, we conclude that NWs based on Pb and Te would be more likely to form this unusual pentagonal structure since they will have weaker ionic bonds.
@@
Further research is necessary to advance towards the alloyed phase Pb1-xSnxTe. Our preliminary results indicate that the difference between the binding energy of the cubic and pentagonal NWs decreases to smaller values in the alloy. Exploring the possibility that the CC consists solely of either Sn or Pb atoms could be a direction worth investigating. In this regard, employing genetic algorithms might predict the crystal structure and the cationic composition of the CCTarkowski and Gonzalez Szwacki (2023).